Out of 525 people surveyed, 210 answered yes. Estimate how many people would answer yes if 900 people were surveyed.
Answer by azhou4(12)

For this one, you would do a proportion. 210/525=x/900
Then you would do cross products. 525x=210*90
You would do 210*90/525. Hope this helps!